# Associative memory and dead neurons

### v1

@@ 1 Introduction
The model in [17] is based on the nonlinear dynamical system that evolves in time from a given initial state. Nonlinear dynamical systems show an exceptionally diverse set of behavior [29], so one needs to select an appropriate class of ordinary differential equations suitable to model associative memory. The most cluical requirement is the ability of the system to evolve to a single state from many initial conditions that are close according to some problem-specific metrics. This fact suggests one should use a dynamical system with many stable steady states. If one selects such a system, each stable steady state corresponds to a particular memory and basin of attraction – all initial conditions that evolve to a selected state – defines a measure of similarity between states.
@@
In this article, we study the Lyapunov function proposed in [17] and later adopted and modified in other recent papers, e.g., [10], [9], [24]. Our main observation is that energy function is vulnerable to the problem of dead neurons [21], which result in a flat energy direction. We illustrate this in the left panel of Figure 1, where one can find both the vector field of the dynamical system and isolines of the energy function. Clearly, the energy function from [17] does not ensure stability and is not helpful in identifying a steady state when dead neurons are present.
@@ 2 Dead neurons
In [17] Krotov and Hopfield proposed the following set of equations to model associative memory
@@
$\boldsymbol{g}(\boldsymbol{y})=\frac{\partial L(\boldsymbol{y})}{\partial\boldsymbol{y}},\,\boldsymbol{W}=\boldsymbol{W}^{\top},$ (2)
$\Lambda_{ij}(\boldsymbol{y})=\frac{\partial^{2}L(\boldsymbol{y})}{\partial y_{i}\partial y_{j}},\,\boldsymbol{\Lambda}(\boldsymbol{y})\geq 0,$ (3)
@@
Equation (1) describes a temporal dynamics of feature vector $\boldsymbol{y}$ starting from $\boldsymbol{y}_{0}$. The equation contains weights $\boldsymbol{W}$, bias $\boldsymbol{b}$ and activation function $\boldsymbol{g}(\boldsymbol{y})$. Structural assumptions (2) ensure that the energy function (4) exists, and restriction on the Hessian $\boldsymbol{\Lambda}$ of Lagrange function $L(\boldsymbol{y})$ (3) guarantees that energy (4) is non-increasing on trajectory of dynamical system (1).
Dynamical system (1) has several favorable properties: (i) steady states can be used as memory vectors, (ii) dynamics in the basins of attraction naturally model memory recovery process, (iii) energy function can be used to ensure the existence of steady states and basins of attraction, (iv) memory is related to neural ODEs so can be trained end-to-end, (v) with a special choice of $\boldsymbol{W}$ steady states resemble activation pattern of classical deep learning architectures. The example below illustrates the last point.
@@
$\dot{\boldsymbol{y}}_{i}=\boldsymbol{W}_{ii-1}\boldsymbol{g}\left(\boldsymbol{y}_{i-1}\right)+\boldsymbol{W}_{ii+1}\boldsymbol{g}\left(\boldsymbol{y}_{i+1}\right)-\boldsymbol{y}_{i}+\boldsymbol{b}_{i},$
@@
Besides MLP described in Example 1 associative memory (1) leads to many more interesting and fruitful connections. In [17] it allowed authors to reconstruct dense associative memory [18] and modern Hopfield network [27]. In [15] and [9] it was used to build a memory model with dense hidden layers and convolution neural networks with pooling layers. In [10] authors introduced an energy transformer using the same formalism. Authors of [30] also noticed that MLP-mixer [31] is related to associative memory (1). Clearly, the technique is valuable and versatile.
@@
For sigmoid activation function $\sigma(x)=\left(1+\exp(-x)\right)^{-1}$ Lagrange function reads $L(\boldsymbol{y})=\sum_{i}\log\left(1+e^{y_{i}}\right)$, and the energy is $E(\boldsymbol{y})=\left(\boldsymbol{y}-\boldsymbol{b}\right)^{\top}\boldsymbol{\sigma}(\boldsymbol{y})-\sum_{i}\log\left(1+e^{y_{i}}\right)-\frac{1}{2}\left(\boldsymbol{\sigma}(\boldsymbol{y})\right)^{\top}\boldsymbol{W}\,\boldsymbol{\sigma}(\boldsymbol{y})$. Neuron $i$ dies when sigmoid saturates for some $y_{i}\gg 1$. After that for all $\widetilde{\boldsymbol{y}}=\boldsymbol{y}+\alpha\boldsymbol{e}_{i}$, with $\alpha\geq 0$ the energy has flat region, i.e., $E(\boldsymbol{y})=E(\widetilde{\boldsymbol{y}}(\alpha))$ for all admissible $\alpha$.
@@
Activation $\text{softmax}\left(\boldsymbol{y}\right)=\exp(\boldsymbol{y})\big{/}\sum_{i=1}^{N}\exp(y_{i})$ corresponds to $L(\boldsymbol{y})=\log\left(\sum_{i=1}^{N}\exp(y_{i})\right)$ and energy $E(\boldsymbol{y})=\left(\boldsymbol{y}-\boldsymbol{b}\right)^{\top}\text{softmax}\left(\boldsymbol{y}\right)-\log\left(\sum_{i=1}^{N}\exp(y_{i})\right)-\frac{1}{2}\left(\text{softmax}\left(\boldsymbol{y}\right)\right)^{\top}\boldsymbol{W}\,\text{softmax}\left(\boldsymbol{y}\right)$. Consider variable $\widetilde{\boldsymbol{y}}=\boldsymbol{y}+\boldsymbol{c}$, where $\boldsymbol{c}$ is a vector with all components equal $c\in\mathbb{R}$. So constant shift, $\text{softmax}\left(\boldsymbol{y}\right)=\text{softmax}\left(\widetilde{\boldsymbol{y}}\right)$, Lagrange function shifts on $c$ and $\widetilde{\boldsymbol{y}}^{\top}\text{softmax}(\widetilde{\boldsymbol{y}})=\boldsymbol{y}^{\top}\text{softmax}(\boldsymbol{y})+c$, so overall energy remains the same, i.e., $E(\boldsymbol{y})=E(\widetilde{\boldsymbol{y}}(c))$ for arbitrary $c$.
The examples above demonstrate that energy function (4) fails to distinguish states on a large fraction of state space. In Examples 2 and 3 this happens because the activation function saturates for some inputs, and in Example 4 this is a consequence of invariance. The latter case is typically not considered as a dead neuron, but since the number of degrees of freedom decreases by one we call this “effective neuron” dead all the same. The formal definition of dead neurons appears below (see [23], [21], [4] for the discussion of dead neurons in the context of ReLU activation function).
@@
The examples above hint that most of the activation functions will lead to dead neurons. The first large class is activation functions with saturation: hyperbolic tangent, sigmoid, ReLU, GELU, SiLU, SELU, Gaussian, etc. The second class is activation functions with symmetries, e.g., softmax and layer norm.
With the definition of dead neurons, we can formalize a pathological behavior of energy function (4).
@@
If at a given point $\boldsymbol{y}$ activation function $\boldsymbol{g}$ has $k$ dead neuron defined by $\boldsymbol{V}\in\mathbb{R}^{N\times k}$, energy function (4) has constant value in a subspace $\mathcal{D}=\left\{\boldsymbol{y}+\boldsymbol{V}c:\boldsymbol{c}\in\mathbb{R}^{k}_{+}\right\}$.
Proof.
Taking $\widetilde{\boldsymbol{y}}=\boldsymbol{y}+\boldsymbol{V}\boldsymbol{c}$, we find from Taylor’s theorem with mean-value form of reminder $L(\widetilde{\boldsymbol{y}})=L(\boldsymbol{y})+\boldsymbol{g}(\boldsymbol{y}+\tau\boldsymbol{V}\boldsymbol{c})^{\top}\boldsymbol{V}\boldsymbol{c}$ for some $\tau\in(0,1)$. Since $\boldsymbol{g}(\boldsymbol{y}+\tau\boldsymbol{V}\boldsymbol{c})=\boldsymbol{g}(\boldsymbol{y})$ by assumption, the second term in the energy function (4) shifts by $-\boldsymbol{g}(\boldsymbol{y})^{\top}\boldsymbol{V}\boldsymbol{c}$ which compensates similar change in the first term, so overall energy does not change under the transformation $\boldsymbol{y}\rightarrow\boldsymbol{y}+\boldsymbol{V}\boldsymbol{c}$. ∎
Flat energy directions are illustrated in Figure 1. Intuitively, it is clear that having large regions with flat energy is not good. In the next section, we explain in detail what will go awry.
@@ 3.1 Sensitivity
and dynamical system is precisely the same as (1). Activation functions $g_{i}$ are invertible and monotone by assumption. The first term of (5) is seemingly distinct from any term of (4). To show that this is an illusion we simplify integrals as follows
@@
    Suppose at a given point $\boldsymbol{y}$ dead neurons are described by matrix $\boldsymbol{V}\in\mathbb{R}^{N\times k}$. In this case energy function $E(\boldsymbol{y},\widetilde{\boldsymbol{b}})$ is Lyapunov function for dynamical systems (1) with $\boldsymbol{b}=\widetilde{\boldsymbol{b}}-\boldsymbol{V}\boldsymbol{c}$ for any $\boldsymbol{c}\in\mathbb{R}^{k}_{+}$.
@@ 3.2 Stability of dynamical system
$E(\boldsymbol{y}^{\star}+\boldsymbol{\delta})-E(\boldsymbol{y}^{\star})=\boldsymbol{\delta}^{\top}\left.\frac{\partial E(\boldsymbol{u})}{\partial\boldsymbol{u}}\right|_{\boldsymbol{u}=\boldsymbol{y}^{\star}}+\boldsymbol{\delta}^{\top}\left.\frac{\partial^{2}E(\boldsymbol{u})}{\partial\boldsymbol{u}^{2}}\right|_{\boldsymbol{u}=\boldsymbol{y}^{\star}+t\boldsymbol{\delta}}\boldsymbol{\delta},\,t(\boldsymbol{\delta},\boldsymbol{y}^{\star})\in(0,1),$
@@
$E(\boldsymbol{y}^{\star}+\boldsymbol{\delta})-E(\boldsymbol{y}^{\star})=\boldsymbol{\delta}^{\top}\left(\boldsymbol{\Lambda}(\boldsymbol{y}^{\star})-\boldsymbol{\Lambda}(\boldsymbol{y}^{\star})\boldsymbol{W}\boldsymbol{\Lambda}(\boldsymbol{y}^{\star})\right)\boldsymbol{\delta},\left\|\boldsymbol{\delta}\right\|_{2}\ll 1$ (6)
For stability in the vicinity of a steady state, it is sufficient that the matrix above is positive definite, and for instability, it is sufficient that at least one eigenvalue of this matrix is negative. In situations such as in Figure 1 when energy has a flat direction nothing can be said about the dynamics in the region with flat energy. It is easy to show that if dead neurons are present at point $\boldsymbol{u}$, matrix $\boldsymbol{V}$ lays in zero eigenspace of $\boldsymbol{\Lambda}(\boldsymbol{u})$. To show this we consider the Taylor series
$\boldsymbol{g}(\boldsymbol{u}+\boldsymbol{V}\boldsymbol{c})=\boldsymbol{g}(\boldsymbol{u})+\boldsymbol{\Lambda}(\boldsymbol{u}+t\boldsymbol{V}\boldsymbol{c})\boldsymbol{V}\boldsymbol{c},\,t\in(0,1),$ (7)
and since $\boldsymbol{g}(\boldsymbol{u}+\boldsymbol{V}\boldsymbol{c})=\boldsymbol{g}(\boldsymbol{u})$ we confirm that $\boldsymbol{\Lambda}\boldsymbol{V}=0$, i.e., zero eigenspace is always present [^1] In the case when energy is flat one may hope to use LaSalle’s invariance principle [7, Theorem 3.3]. Unfortunately, it is also not applicable, because flat regions are not compact, so the existence of a positive invariant set can not be ensured.
@@
where $\boldsymbol{P}_{d}=\boldsymbol{V}\boldsymbol{V}^{\top}$ and $\boldsymbol{P}_{a}=\boldsymbol{I}-\boldsymbol{P}_{d}$. One can immediately see that dead neurons do not influence stability and, when steady state $\boldsymbol{y}_{a}^{\star}$ is reached, one can find $\boldsymbol{y}_{d}^{\star}=\boldsymbol{P}_{d}\boldsymbol{W}\boldsymbol{g}(\boldsymbol{y}_{a}^{\star})+\boldsymbol{P}_{d}\boldsymbol{b}$. Moreover, when $\boldsymbol{y}_{a}(t)$ reaches steady state $\boldsymbol{y}_{d}(t)$ converges exponentially from the arbitrary starting point, i.e., the whole flat region is a basin of attraction.
@@
The structure of matrix $\boldsymbol{S}$ suggests that, in fact, dynamic is stable when energy function indicates that. Indeed, $\boldsymbol{\Lambda}_{\perp}^{1/2}\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}^{1/2}$ has the same spectrum as $\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}$, so sufficient condition for $\boldsymbol{S}$ is equivalent to sufficient condition for $\boldsymbol{\Lambda}_{\perp}^{1/2}\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}^{1/2}-\boldsymbol{I}$. Finally since $\boldsymbol{\Lambda}_{\perp}$ is full rank we find that if $\boldsymbol{\Lambda}_{\perp}\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}-\boldsymbol{\Lambda}_{\perp}$ is negatively stable (has non-positive eigenvalues), the dynamics of $\boldsymbol{y}_{a}$ is stable. Since this matrix is a restriction of $\boldsymbol{\Lambda}\boldsymbol{W}\boldsymbol{\Lambda}-\boldsymbol{\Lambda}$ on the range of Hessian we conclude that one can analyze the stability of steady-state using energy if projector on the range of Hessian is known. It is important to note that matrix $\boldsymbol{\Lambda}\boldsymbol{W}\boldsymbol{\Lambda}-\boldsymbol{\Lambda}$ can have other flat directions not corresponding to nullspace of Hessian, these directions will cause instability, so one need to distinguish them from directions corresponding to dead neurons.
@@
Proof.
Matrix $\boldsymbol{W}_{\perp}$ can be written as $\boldsymbol{D}^{\top}\boldsymbol{O}^{\top}\boldsymbol{W}\boldsymbol{O}\boldsymbol{D}$ where $\boldsymbol{O}$ is orthogonal full rank matrix and $\boldsymbol{D}$ is first $N-k$ columns of $\boldsymbol{I}$. Since GOE is invariant under orthogonal transformations $\boldsymbol{O}^{\top}\boldsymbol{W}\boldsymbol{O}$ is also a matrix from GOE. The effect of $\boldsymbol{D}$ is to select a principal submatrix with $N-k$ rows and columns. Given that $\boldsymbol{W}_{\perp}$ is also from GOE but with $N-k$ rows and columns. For GOE with large $N$, the distribution of leading eigenvalue quickly converges to Tracy–Widom distribution (see [3] equation (50) and numerical results), so, on the average, spectral radius of a matrix with $N-k$ rows and columns is $\sqrt{2(N-k)}$ and $\mathbb{E}\left\|\boldsymbol{W}_{\perp}\right\|_{2}\big{/}\mathbb{E}\left\|\boldsymbol{W}\right\|_{2}=\sqrt{1-\frac{k}{N}}$. ∎

### v2

@@ 1 Introduction
The model in [17] is based on the nonlinear dynamical system that evolves in time from a given initial state. Nonlinear dynamical systems show an exceptionally diverse set of behavior [29], so one needs to select an appropriate class of ordinary differential equations suitable to model associative memory. The most crucial requirement is the ability of the system to evolve to a single state from many initial conditions that are close according to some problem-specific metrics. This fact suggests one should use a dynamical system with many stable steady states. If one selects such a system, each stable steady state corresponds to a particular memory and basin of attraction – all initial conditions that evolve to a selected state – defines a measure of similarity between states.
@@
In this article, we study the Lyapunov function proposed in [17] and later adopted and modified in other recent papers, e.g., [10], [9], [24]. Our main observation is that energy function is vulnerable to the problem of dead neurons [^1] Usually, neurons are called dead when activation function saturates, e.g., $x<0$ for ReLU or $x\gg 1$ for sigmoid. In this article we use an extended definition of dead neurons and say that a neuron is dead when it is impossible to reconstruct input to activation function from its output. This is explained in more detail in Section 2. [21], which result in a flat energy direction. We illustrate this in the left panel of Figure 1, where one can find both the vector field of the dynamical system and isolines of the energy function. Clearly, the energy function from [17] does not ensure stability and is not helpful in identifying a steady state when dead neurons are present.
@@ 2 Dead neurons
In [17] authors proposed to rewrite activation function $\boldsymbol{g}(\boldsymbol{y})$ as gradient of Lagrange function $L(\boldsymbol{y})$, i.e., $\boldsymbol{g}(\boldsymbol{y})=\frac{\partial L(\boldsymbol{y})}{\partial\boldsymbol{y}}$. This formalism allows to describe a large class of memory models with Lyapunov function on the common grounds. Results that we obtain in the article are applicable to all of them, but for simplicity we consider dense model taken from [15, Equation (2)] (see Appendix A or further discussion on relations between different models). This model reads:
@@
Equation (1) describes a temporal dynamics of feature vector $\boldsymbol{y}$ starting from $\boldsymbol{y}_{0}$. The equation contains weights $\boldsymbol{W}$, bias $\boldsymbol{b}$ and activation function $\boldsymbol{g}(\boldsymbol{y})$. Energy function (2) is non-increasing on trajectories of (1) if and only if $\boldsymbol{W}=\boldsymbol{W}^{\top}$ and Hessian of Lagrange function $\boldsymbol{\Lambda}=\frac{\partial^{2}L(\boldsymbol{y})}{\partial\boldsymbol{y}^{2}}$ is positive semi-definite (see [17, Equation (4), Appendix A] and [15, Equation (4), Appendix A]).
Dynamical system (1) has several favorable properties (see discussion in Section 1 and [17] , [15]): (i) steady states can be used as memory vectors, (ii) dynamics in the basins of attraction naturally model memory recovery process, (iii) energy function can be used to ensure the existence of steady states and basins of attraction, (iv) memory is related to neural ODEs so can be trained end-to-end, (v) with a special choice of $\boldsymbol{W}$ steady states resemble activation pattern of classical deep learning architectures. The example below illustrates the last point (see [15, Section 3]).
@@
$\begin{split}&\dot{\boldsymbol{y}}_{i}=\boldsymbol{W}_{i,i-1}\boldsymbol{g}\left(\boldsymbol{y}_{i-1}\right)+\boldsymbol{W}_{i,i+1}\boldsymbol{g}\left(\boldsymbol{y}_{i+1}\right)-\boldsymbol{y}_{i}+\boldsymbol{b}_{i},\,{i=2,3}\\ &{\dot{\boldsymbol{y}}_{1}=\boldsymbol{W}_{12}\boldsymbol{g}\left(\boldsymbol{y}_{2}\right)-\boldsymbol{y}_{1}+\boldsymbol{b}_{1},\,\dot{\boldsymbol{y}}_{4}=\boldsymbol{W}_{43}\boldsymbol{g}\left(\boldsymbol{y}_{3}\right)-\boldsymbol{y}_{4}+\boldsymbol{b}_{4},}\end{split}$
@@
Besides MLP described in Example 1 associative memory (1) leads to many more interesting and fruitful connections. In [17] it allowed the authors to reconstruct dense associative memory [18] and modern Hopfield network [27]. In [15] and [9] it was used to build a memory model with dense hidden layers and convolution neural networks with pooling layers. In [10] authors introduced an energy transformer using the same formalism. The authors of [30] also noticed that MLP-mixer [31] is related to associative memory (1). Clearly, the technique is valuable and versatile.
@@
For sigmoid activation function $\sigma(x)=\left(1+\exp(-x)\right)^{-1}$ Lagrange function reads $L(\boldsymbol{y})=\sum_{i}\log\left(1+e^{y_{i}}\right)$, and the energy is $E(\boldsymbol{y})=\left(\boldsymbol{y}-\boldsymbol{b}\right)^{\top}\boldsymbol{\sigma}(\boldsymbol{y})-\sum_{i}\log\left(1+e^{y_{i}}\right)-\frac{1}{2}\left(\boldsymbol{\sigma}(\boldsymbol{y})\right)^{\top}\boldsymbol{W}\,\boldsymbol{\sigma}(\boldsymbol{y})$. Neuron $i$ dies when sigmoid saturates222 Formally, the sigmoid function saturates in the limit $y\rightarrow\pm\infty$. However, for computations in floating-point arithmetic it is safe to say that sigmoid saturates when the absolute value of the argument is sufficiently large but finite. for some $y_{i}\gg 1$. After that for all $\widetilde{\boldsymbol{y}}=\boldsymbol{y}+\alpha\boldsymbol{e}_{i}$, with $\alpha\geq 0$ the energy has flat region, i.e., $E(\boldsymbol{y})=E(\widetilde{\boldsymbol{y}}(\alpha))$ for all admissible $\alpha$.
@@
Activation $\text{softmax}\left(\boldsymbol{y}\right)=\exp(\boldsymbol{y})\big{/}\sum_{i=1}^{N}\exp(y_{i})$ corresponds to $L(\boldsymbol{y})=\log\left(\sum_{i=1}^{N}\exp(y_{i})\right)$ and energy $E(\boldsymbol{y})=\left(\boldsymbol{y}-\boldsymbol{b}\right)^{\top}\text{softmax}\left(\boldsymbol{y}\right)-\log\left(\sum_{i=1}^{N}\exp(y_{i})\right)-\frac{1}{2}\left(\text{softmax}\left(\boldsymbol{y}\right)\right)^{\top}\boldsymbol{W}\,\text{softmax}\left(\boldsymbol{y}\right)$. Consider variable $\widetilde{\boldsymbol{y}}{(\boldsymbol{c})}=\boldsymbol{y}+\boldsymbol{c}$, where $\boldsymbol{c}$ is a vector with all components equal $c\in\mathbb{R}$. So constant shift, $\text{softmax}\left(\boldsymbol{y}\right)=\text{softmax}\left(\widetilde{\boldsymbol{y}}{(\boldsymbol{c})}\right)$, Lagrange function shifts on $c$ and $\widetilde{\boldsymbol{y}}{(\boldsymbol{c})}^{\top}\text{softmax}(\widetilde{\boldsymbol{y}}{(\boldsymbol{c})})=\boldsymbol{y}^{\top}\text{softmax}(\boldsymbol{y})+c$, so overall energy remains the same, i.e., $E(\boldsymbol{y})=E(\widetilde{\boldsymbol{y}}(c))$ for arbitrary $c$.
The examples above demonstrate that energy function (2) fails to distinguish states on a large fraction of state space. In Examples 2 and 3 this happens because the activation function saturates for some inputs, and in Example 4 this is a consequence of invariance. The latter case is typically not considered as a dead neuron, but since the number of degrees of freedom decreases by one we call this “effective neuron” dead all the same. [^3] One can also say that softmax is always saturated. If we consider input $\boldsymbol{y}$ in the new basis where $\widetilde{y}_{1}=\sum_{i}y_{i}$, it is easy to see that after the softmax this component becomes $1$ regardless of the input. The formal definition of dead neurons appears below (see [23], [21], [4] for the discussion of dead neurons in the context of ReLU activation function).
@@
The examples above hint that most of the activation functions will lead to dead neurons. The first large class is activation functions with saturation: hyperbolic tangent, sigmoid, ReLU, GELU, SiLU, SELU, Gaussian, etc.444 For sigmoid, hyperbolic tangent and Gaussian function the saturation is precise only in the limit $x\rightarrow\pm\infty$. However, as explained in the Example 3 in the floating-point arithmetics these functions effectively saturate for finite values of the argument. Moreover, even with arbitrary precision, approximately flat direction can not be used to reliably store any information since it will require weights with exponentially large magnitudes. The second class is activation functions with symmetries, e.g., softmax and layer norm. With the definition of dead neurons, we can formalize a pathological behavior of energy function (2).
@@
If at a given point $\boldsymbol{y}$ activation function $\boldsymbol{g}$ has $k$ dead neuron defined by $\boldsymbol{V}\in\mathbb{R}^{N\times k}$, energy function (2) has constant value in a subspace $\mathcal{D}=\left\{\boldsymbol{y}+\boldsymbol{V}{\boldsymbol{c}}:\boldsymbol{c}\in\mathbb{R}^{k}_{+}\right\}$.
Proof: Appendix B.
Flat energy directions are illustrated in Figure 1 and Figure 2(c). Intuitively, it is clear that having large regions with flat energy is not good. In the next section, we explain in detail what will go awry.
As illustrated in Figure 2 energy function can appear problematic for reasons not directly related to dead neurons. We argue that these two situations are less grim. Energy illustrated in Figure 2(a) can still be utilized with proper initial conditions, e.g., $\left\|\boldsymbol{y}_{0}\right\|_{2}\leq R$ for some small $R$. Besides, we demonstrate in Appendix C that restricting parameters to ensure energy is bounded from below can lead to catastrophic capacity reduction. Energy from Figure 2(b) is less problematic for stability analysis because invariance principle [7, Theorem 3.3] can be used.
@@ 3 Consequences of dead neurons
Figure 2: Sketch of three problematic energy functions: (a) unbounded from below with two stable states (we analyze this situation in Appendix C), (b) bounded from below but with compact flat region, (c) bounded from below but with non-compact flat region. According to Proposition 1 case (c) is realized in models [17] when neurons die. In red regions stability properties do not follow from Lyapunov theorems. LaSalle’s invariance principle [7, Theorem 3.3] ensures that in case (b) isolated steady states are stable, but for the case (c) separate stability analysis is needed (see Section 3.2 for details).
@@ 3.1 Sensitivity
and dynamical system is precisely the same as (1). Activation functions $g_{i}$ are invertible and monotone by assumption. In Appendix D we explain precisely how to adapt the model from [13] to our context. The first term of (3) is seemingly distinct from any term of (2). To show that this is an illusion we simplify integrals as follows
@@
    Suppose at a given point $\boldsymbol{y}$ dead neurons are described by matrix $\boldsymbol{V}\in\mathbb{R}^{N\times k}$. In this case energy function $E(\boldsymbol{y})$ is Lyapunov function for dynamical systems (1) with $\boldsymbol{b}=\widetilde{\boldsymbol{b}}-\boldsymbol{V}\boldsymbol{c}$ for any $\boldsymbol{c}\in\mathbb{R}^{k}_{+}$.
@@ 3.2 Stability of dynamical system
$E(\boldsymbol{y}^{\star}+\boldsymbol{\delta})-E(\boldsymbol{y}^{\star})=\boldsymbol{\delta}^{\top}\left.\frac{\partial E(\boldsymbol{u})}{\partial\boldsymbol{u}}\right|_{\boldsymbol{u}=\boldsymbol{y}^{\star}}+{\frac{1}{2}}\boldsymbol{\delta}^{\top}\left.\frac{\partial^{2}E(\boldsymbol{u})}{\partial\boldsymbol{u}^{2}}\right|_{\boldsymbol{u}=\boldsymbol{y}^{\star}+t\boldsymbol{\delta}}\boldsymbol{\delta},\,t(\boldsymbol{\delta},\boldsymbol{y}^{\star})\in(0,1),$
@@
$E(\boldsymbol{y}^{\star}+\boldsymbol{\delta})-E(\boldsymbol{y}^{\star})=\boldsymbol{\delta}^{\top}\left(\boldsymbol{\Lambda}(\boldsymbol{y}^{\star})-\boldsymbol{\Lambda}(\boldsymbol{y}^{\star})\boldsymbol{W}\boldsymbol{\Lambda}(\boldsymbol{y}^{\star})\right)\boldsymbol{\delta},\left\|\boldsymbol{\delta}\right\|_{2}\ll 1.$
For stability in the vicinity of a steady state, it is sufficient that the matrix above is positive definite, and for instability, it is sufficient that at least one eigenvalue of this matrix is negative. In situations such as in Figure 1 and Figure 2(c) when energy has a flat direction nothing can be said about the dynamics in this region [^5] Note, however, that for the face given in Figure 2(b) it is possible to construct a positively invariant set based on energy levels. After that LaSalle’s invariance principle [7, Theorem 3.3] guarantees that isolated steady states are asymptotically stable. This material is standard and covered in many books, e.g., [7, Section 3.3.].. It is easy to show that if dead neurons are present at point $\boldsymbol{u}$, matrix $\boldsymbol{V}$ lays in zero eigenspace of $\boldsymbol{\Lambda}(\boldsymbol{u})$. To show this we consider the Taylor series $\boldsymbol{g}(\boldsymbol{u}+\boldsymbol{V}\boldsymbol{c})=\boldsymbol{g}(\boldsymbol{u})+\boldsymbol{\Lambda}(\boldsymbol{u}+t\boldsymbol{V}\boldsymbol{c})\boldsymbol{V}\boldsymbol{c},\,t\in(0,1),$ and since $\boldsymbol{g}(\boldsymbol{u}+\boldsymbol{V}\boldsymbol{c})=\boldsymbol{g}(\boldsymbol{u})$ we confirm that $\boldsymbol{\Lambda}\boldsymbol{V}=0$, i.e., zero eigenspace is always present.
@@
where $\boldsymbol{P}_{d}=\boldsymbol{V}\boldsymbol{V}^{\top}$ and $\boldsymbol{P}_{a}=\boldsymbol{I}-\boldsymbol{P}_{d}$. [^6] We assume for simplicity that $\boldsymbol{V}$ has orthonormal columns. One can immediately see that dead neurons do not influence stability and, when steady state $\boldsymbol{y}_{a}^{\star}$ is reached, one can find $\boldsymbol{y}_{d}^{\star}=\boldsymbol{P}_{d}\boldsymbol{W}\boldsymbol{g}(\boldsymbol{y}_{a}^{\star})+\boldsymbol{P}_{d}\boldsymbol{b}$. Moreover, when $\boldsymbol{y}_{a}(t)$ reaches steady state $\boldsymbol{y}_{d}(t)$ converges exponentially from the arbitrary starting point, i.e., the whole flat region is a basin of attraction.
@@
The structure of matrix $\boldsymbol{S}$ suggests that, in fact, dynamics is stable when energy function indicates that. Indeed, $\boldsymbol{\Lambda}_{\perp}^{1/2}\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}^{1/2}$ has the same spectrum as $\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}$, so sufficient condition for $\boldsymbol{S}$ is equivalent to sufficient condition for $\boldsymbol{\Lambda}_{\perp}^{1/2}\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}^{1/2}-\boldsymbol{I}$. Finally since $\boldsymbol{\Lambda}_{\perp}$ is full rank we find that if $\boldsymbol{\Lambda}_{\perp}\boldsymbol{W}_{\perp}\boldsymbol{\Lambda}_{\perp}-\boldsymbol{\Lambda}_{\perp}$ is negatively stable (has non-positive eigenvalues), the dynamics of $\boldsymbol{y}_{a}$ is stable. Since this matrix is a restriction of $\boldsymbol{\Lambda}\boldsymbol{W}\boldsymbol{\Lambda}-\boldsymbol{\Lambda}$ on the range of Hessian we conclude that one can analyze the stability of steady-state using energy if projector on the range of Hessian is known. It is important to note that matrix $\boldsymbol{\Lambda}\boldsymbol{W}\boldsymbol{\Lambda}-\boldsymbol{\Lambda}$ can have other flat directions not corresponding to nullspace of Hessian, these directions will cause instability, so one need to distinguish them from directions corresponding to dead neurons.
@@
Proof: Appendix F.